OpenStudy (poopsiedoodle):

1. A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air? 2. Two bicycles depart from Miami Beach going in opposite directions. The first bicycle is traveling at 10 miles per hour. The second bicycle travels at 5 miles per hour. How long does it take until the bikes are 45 miles apart?

OpenStudy (poopsiedoodle):

3. Jesse rents a moving van for $75 and must pay $2 per mile. The following week, Alex rents the same van, is charged $80 for the rental and $1.50 per mile. If they each paid the same amount and drove the same number of miles, how far did they each travel? 4. During a 4th of July weekend, 32 vehicles became trapped on the Sunshine Skyway Bridge while it was being repaved. A recent city ordinance decreed that only cars with 4 wheels and trucks with six wheels could be on the bridge at any given time. If there were 148 tires that needed to be replaced to due to damage, how many cars and trucks were involved in the incident?

OpenStudy (anonymous):

How any questions are there?

OpenStudy (poopsiedoodle):

Just these and one more where I have to do something with another person.

OpenStudy (anonymous):

You should post these one by one. It makes it less confusing

OpenStudy (poopsiedoodle):

but they're numbered...

OpenStudy (anonymous):

Hint with question one since the 40mph tailwind acts for the same duration, in opposite directions it cancels itself out. for this your really just working out how much speed a plane has when it travels 5 miles in six hours \[distance = \frac{velocity}{time}\]

OpenStudy (poopsiedoodle):

Stilllllllll clueless.

OpenStudy (anonymous):

distance = speed / time. we want to know speed. so if we solve the equation for speed then put in the values of distance, 5 miles, and time, 6 hours, we get the speed.

OpenStudy (anonymous):

by we i mean you :P

OpenStudy (poopsiedoodle):

It's 5 hours, not 5 miles o.o

OpenStudy (anonymous):

ಠ_ಠ

OpenStudy (poopsiedoodle):

ლ(ಠ益ಠლ) Y U NO ANSWER?

OpenStudy (anonymous):

okay wel swap them then! what is speed measured in? meters per second. what is distance? meters. time? seconds. So we have\[meters per second = \frac{meters}{second} = \frac{distance}{time}\] right?

OpenStudy (poopsiedoodle):

miles per hour*

OpenStudy (anonymous):

lol im hopeless at explaining

OpenStudy (poopsiedoodle):

I got 10 for #3.

OpenStudy (poopsiedoodle):

and I got 22 cars and 10 trucks for #4